OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. 1 E(-2, 2), F(5, 1) 2 G(4, 6), H(1, 5) 3 L(5, 0), M(0, 5)

OpenStudy (mertsj):

The slope of the first one is 1/-7 y-1=-1/7(x-5) 1/7x+y=12/7 or x+7y=12

OpenStudy (anonymous):

thanks what about the others

OpenStudy (mertsj):

Have you tried either of them?

OpenStudy (anonymous):

yeah but i cant qet it

OpenStudy (mertsj):

Slope of second one is 1/3 y-5=1/3(x-1) -1/3x+y=14/3 or x-3y=-14

OpenStudy (mertsj):

Third one slope is -1 y-5=-1(x-0) x+y=5

OpenStudy (anonymous):

thanks man alot